The bird in fig A drops the stone from a height of 88 m above the water surface. Initially, the stone has zero vertical velocity. How long will it take the stone to reach the surface of the pond? (Assume air resistance is negligible.)


1
Expert's answer
2021-01-05T12:04:39-0500

The distance travelled during the free fall is given as follows:


"d = \dfrac{gt^2}{2}"

where "g = 9.81m\/s^2" is the gravitational acceleration, "t" is the time that was taken the stone to reach the surface of the pond. Since "d = 88m", obtain:


"t = \sqrt{\dfrac{2d}{g}}\\
t = \sqrt{\dfrac{2\cdot 88}{9.81}} \approx 4.2s"

Answer. 4.2 s.